# A JWST Project on 47 Tucanae: Kinematics, energy equipartition and anisotropy of multiple populations

### v1

@@
22institutetext: Istituto Nazionale di Astrofisica - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova, IT-35122 33institutetext: Research School of Astronomy & Astrophysics, Australian National University, Canberra, ACT 2611, Australia 44institutetext: Department of Astrophysics, University of Vienna, Türkenschanzstrasse 17, 1180 Vienna, Austria 55institutetext: Department of Astronomy, Indiana University, Bloomington, IN 47401, USA 66institutetext: Department of Physics and Astronomy, Sejong University, 209 Neungdong-ro, Gwangjin-Gu, Seoul, 05006, Republic of Korea 77institutetext: Space Telescope Science Institute, 3700 San Martin Dr, Baltimore, MD, 21218, USA 88institutetext: South-Western Institute for Astronomy Research, Yunnan University, Kunming 650500, P. R. China 99institutetext: Universidade Federal do Rio Grande do Sul, Instituto de Física, Av. Bento Gonçalves 9500, Porto Alegre, RS, Brazil
@@ A JWST Project on 47 Tucanae: Kinematics, energy equipartition and anisotropy of multiple populations
T. Ziliotto 11    A. P. Milone 1122    G. Cordoni 33    F. I. Aros 44    E. Vesperini 55    J. W. Lee 66    A. Bellini 77    M. Libralato 22    E. Dondoglio 11    M. Tailo 11    A. Livernois 33    M. V. Legnardi 11    A. Mastrobuono-Battisti 11    E. Lagioia 99    E. Bortolan 11    F. Muratore 11    A. F. Marino 22    A. Alves-Brito 1010    A. Renzini 22
@@ 1 Introduction
The origin of this phenomenon has been widely discussed, and multiple scenarios have been proposed to explain the presence of multiple populations (Cottrell & Da Costa 1981; Dantona et al. 1983; Decressin et al. 2007; Bastian et al. 2013; Renzini et al. 2015; D’Antona et al. 2016; D’Ercole et al. 2016; Calura et al. 2019; Wang et al. 2020; Renzini et al. 2022). Recent observational results with data from the James Webb Space Telescope (JWST) and the Hubble Space Telescope (HST) appear to support models predicting multiple episodes of star formation within the same cluster (Ziliotto et al. 2023; Milone et al. 2023; Marino et al. 2024a, b).
A number of studies based on hydro/N-body simulations have investigated the formation of multiple generations in globular clusters and predicted that 2G stars form centrally concentrated in the inner regions of the 1G system (see e.g. D’Ercole et al. 2008; Bekki 2011; Bekki et al. 2017; Calura et al. 2019; Lacchin et al. 2021, 2022; Yaghoobi et al. 2022a, b). While long-term dynamical evolution can gradually erase information of the initial conditions of these systems, some clusters can still preserve this information (see e.g. Vesperini et al. 2021; Livernois et al. 2024), as numerous observational studies demonstrate (Bellini et al. 2015, 2018; Milone et al. 2018; Dalessandro et al. 2019, 2024; Libralato et al. 2023; Leitinger et al. 2023; Cordoni et al. 2020, 2024; Mehta et al. 2024).
GCs are collisional systems, and as their stars interact and exchange energy they evolve toward energy equipartition, a dynamical state where stars with different masses have equal kinetic energy, $K=\frac{1}{2}m\sigma^{2}$, and thus a velocity dispersion, $\sigma$, increasing for smaller stellar masses (see e.g. Spitzer 1969). Theoretical and observational works suggest that GCs do not reach full energy equipartition, but are characterized by a state of partial equipartition (see e.g. Trenti & van der Marel 2013; Bianchini et al. 2016; Libralato et al. 2018; Watkins et al. 2022; Pavlík & Vesperini 2021, 2022; Aros & Vesperini 2023).
@@
However, measuring the degree of energy equipartition of the multiple populations in GCs is challenging. Generally, high-precision proper motion measurements in GCs are available only for bright stars, such as those on the red giant branch (RGB), asymptotic giant branch (AGB), and horizontal branch (HB), which limits the analysis to a narrow mass range. This excludes the entire main sequence (MS), which encompasses a significantly broader range of stellar masses compared to giant stars. Furthermore, high-precision photometry with filters that effectively separate multiple populations is also required.
@@
In the past few years, thanks to progress in measuring kinematics of low-mass stars with the HST, the measurement of energy equipartition in GCs has become possible. Watkins et al. (2022) measured the degree of energy equipartition in the central regions of nine GCs by using HST photometry, finding that no cluster is at or near full equipartition. Similarly, Libralato et al. (2022) found a partial state of energy equipartition in NGC 5904. Heyl et al. (2017) investigated mass-dependent kinematics in 47 Tucanae, finding a weak dependence of proper motions on stellar mass just beyond the cluster’s half-light radius, and a stronger dependence at the cluster center. In the context of multiple stellar populations, only one cluster has been studied, and over small radial ranges. Bellini et al. (2018) measured the degree of energy equipartition for 1G and 2G stars in $\omega$ Centauri, finding that 2G stars present some degree of energy equipartition, while 1G stars are not in equipartition.
@@
Recently, we initiated a project to investigate the GC 47 Tucanae and its stellar populations using JWST observations from GO 2560 (PI: A. F. Marino) and archival data from major telescopes, including HST, JWST, and Gaia. Early results from this project include: i) the identification and chemical characterization of multiple stellar populations among very-low-mass stars using deep Near Infrared Camera (NIRCam) photometry (Milone et al. 2023; Marino et al. 2024a; Legnardi et al. 2024); ii) the first detection of the brown dwarf sequence and the discovery of gaps and discontinuities within it (Marino et al. 2024a); iii) the inaugural spectroscopic study of multiple populations among M-dwarfs, based on Near-Infrared Spectrograph (NIRSpec) data (Marino et al. 2024b); and iv) the exploration of binary systems among first-generation and second-generation stars (Milone et al. 2024, submitted).
@@
In this paper, we continue our investigation into the multiple stellar populations of 47 Tucanae focusing on the kinematics of these populations. By leveraging a combination of datasets from JWST, HST, Gaia, and ground-based telescopes, we identify stellar populations across a mass range from $\sim$0.1 to 0.9 $M_{\odot}$ and derive precise proper motions for these stars. For the first time, we measure the differences in the degree of energy equipartition between 1G and 2G stars in 47 Tucanae, within an extensive radial range from approximately 1.5 to 4.5 $R_{h}$. Additionally, we analyze the internal kinematics of 1G and 2G stars, including anisotropy profiles, over an extensive radial range that extends from the center of the cluster to over 30 arcminutes. This paper is organized as follows: Section 2 describes our photometric and astrometric catalogs; in Section 3, we outline our procedure for deriving stellar masses; the methodology and results regarding velocity dispersion and anisotropy profiles are discussed in Section 4; Section 5 presents the energy equipartition models and results; and Section 6 provides a discussion of the results, including energy equipartition radial profiles for 1G and 2G stars.
@@ 2 Data
In this section, we describe the dataset used to identify multiple stellar populations in the various fields of view analyzed in this study, illustrated in Figure 1. The cumulative exposure times across different years of observation are illustrated in the histogram in Figure 2.
@@
Field A, situated approximately 7 arcminutes from the cluster center, has been extensively observed with the HST, providing deep optical and near-infrared data. With this dataset, we were able to identify multiple populations among lower- ($M<0.45M_{\odot}$) and upper-MS ($0.780.67<M/M_{\odot}<$ 0.78) stars. For a more comprehensive description of this dataset, see Milone et al. (2023).
Fields B and C, situated at $\sim 8.5$ and $\sim 11$ arcminutes from the cluster center, respectively, were observed with JWST/NIRCam as part of the program GO-2560 (PI: A. F. Marino). To classify populations in these fields, we combined NIRCam photometry in the F115W and F322W2 filters with observations taken with the HST Ultraviolet-VISible (UVIS) channel of the Wide Field Camera 3 (WFC3) in the F606W filter, and with NIR/WFC3 in the F110W and F160W filters. See Marino et al. (2024a) and Milone et al. (2023) for a detailed description of these datasets.
@@
To analyze the kinematics of more massive 1G and 2G stars, we used three catalogs: Cordoni et al. (2024), Mehta et al. (2024), and Lee (2022). Mehta et al. (2024) used Gaia DR3 XP spectra (Gaia Collaboration et al. 2023) to derive synthetic photometric filters that distinguish multiple populations in the RGB, HB, and AGB up to the tidal radius of 47 Tucanae. Using ground-based photometry from Stetson et al. (2019), Cordoni et al. (2024) applied the $\Delta_{C_{U,B,I}}$ vs. $\Delta_{B,I}$ chromosome map (ChM) to separate 1G and 2G RGB stars in 47 Tucanae, up to distances of about 20 arcminutes. Finally, Lee (2022) compiled large field-of-view ($\sim 1^{\circ}\times 1^{\circ}$) ground-based photometry in filters designed to maximize the separation of 1G and 2G stars. We matched this catalogs with Gaia DR3 proper motions to examine kinematic differences between 1G and 2G stars. Together, the catalogs from Lee (2022); Mehta et al. (2024); Cordoni et al. (2024) cover stars from about 5 arcminutes to the tidal radius of 47 Tucanae.
@@
Figure 3: Photometric diagrams demonstrating the photometry used to disentangle the multiple populations in different radial regions of 47 Tucanae. Panel a) shows the HST UV photometry used to identify populations in the RGB. In panel b), we show the CMD created using HST optical filters for Field A. Panel c) demonstrates the CMD created for the lower MS and substellar objects with JWST filters for Field C. Panel d) displays HST IR photometry for disentangling populations among the lower MS. In panel e) we show Ca-CN photometry by Lee (2022), which is effective in identifying populations among RGB and HB stars. Panel f) displays the diagram with ground-based UBVI photometry from Stetson et al. (2019). Finally, in panel g), we show the CMD created with Gaia XP spectro-photometry (Mehta et al. 2024).
@@
2.1 Proper Motions
To assess relative proper motions, we employed a methodology detailed in several previous studies (e.g., Anderson & King 2003; Piotto et al. 2012; Libralato et al. 2022; Milone et al. 2023), which focuses on comparing stellar positions across images captured at different epochs.
@@ 2.1 Proper Motions
To transform star coordinates from each catalog into this master frame, we utilized six-parameter linear transformations. To enhance alignment and minimize small residual distortions, accounting for CTE and geometric distortion residuals in the HST (ACS and UVIS) images, we implemented local transformations based on the nearest N reference stars, deliberately excluding target stars themselves from influencing these corrections (as described in Bellini et al. 2014). The selection of N and the corresponding magnitude range was optimized to balance capturing local variations while maintaining a statistically significant sample of reference stars. For stars near the extremes of the magnitude range, either very bright or very faint targets, we adjusted the magnitude selection criteria by using reference stars within the brightest or faintest magnitude bins instead of applying a fixed magnitude difference. This adaptive approach enhanced the available statistics and improved the correction process in regions where suitable reference stars were otherwise scarce.
We then plotted the abscissa and ordinate of each star against the observation epochs, fitting a weighted least-squares linear regression to these data points. The slope of this fitted line, along with its uncertainty, yields our best estimate of the star’s proper motion.
The transformation process relied on bright, unsaturated cluster stars. This selection was conducted in two phases: initially, we identified likely members of 47 Tucanae based solely on their positions in the CMD, which allowed us to derive preliminary proper motions. In the second phase, we refined our criteria by examining each star’s position within both the CMD and the proper motion diagram, leading to a more accurate identification of cluster members. These members were then used to calculate improved proper motions. To further reduce uncertainties in our measurements, we adjusted stellar coordinates at each epoch by accounting for displacements due to proper motions. These corrected coordinates were then utilized to calculate improved transformations, further refining our estimates of proper motion.

### v2

@@
22institutetext: Istituto Nazionale di Astrofisica - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova, IT-35122 33institutetext: Research School of Astronomy & Astrophysics, Australian National University, Canberra, ACT 2611, Australia 44institutetext: Department of Astrophysics, University of Vienna, Türkenschanzstrasse 17, 1180 Vienna, Austria 55institutetext: Department of Astronomy, Indiana University, Bloomington, IN 47401, USA 66institutetext: Department of Physics and Astronomy, Sejong University, 209 Neungdong-ro, Gwangjin-Gu, Seoul, 05006, Republic of Korea 77institutetext: Space Telescope Science Institute, 3700 San Martin Dr, Baltimore, MD, 21218, USA 88institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, F-67000 Strasbourg, France 99institutetext: South-Western Institute for Astronomy Research, Yunnan University, Kunming 650500, P. R. China 1010institutetext: Universidade Federal do Rio Grande do Sul, Instituto de Física, Av. Bento Gonçalves 9500, Porto Alegre, RS, Brazil
@@ A JWST project on 47 Tucanae: kinematics, energy equipartition and anisotropy of multiple populations
T. Ziliotto 11    A. P. Milone 1122    G. Cordoni 33    F. I. Aros 44    E. Vesperini 55    J.-W. Lee 66    A. Bellini 77    P. Bianchini 88    A. Mastrobuono-Battisti 11    M. Libralato 22    E. Dondoglio 22    M. Tailo 11    A. Livernois 55    M. V. Legnardi 11    E. Lagioia 99    E. Bortolan 11    F. Muratore 11    A. F. Marino 22    A. Alves-Brito 1010    A. Renzini 22
@@ 1 Introduction
The origin of this phenomenon has been widely discussed, and multiple scenarios have been proposed to explain the presence of multiple populations (Cottrell & Da Costa 1981; Dantona et al. 1983; Decressin et al. 2007; Bastian et al. 2013; Renzini et al. 2015; D’Antona et al. 2016; D’Ercole et al. 2016; Calura et al. 2019; Wang et al. 2020; Renzini et al. 2022). Recent studies using data from the James Webb Space Telescope (JWST) and the Hubble Space Telescope (HST) have analyzed the chemical variations between 1G and 2G very low-mass stars in M 92 and 47 Tucanae (Ziliotto et al. 2023; Milone et al. 2023b; Marino et al. 2024a, b) and appear to support models predicting multiple episodes of star formation within the same cluster.
A number of studies based on hydro/N-body simulations have investigated the formation of multiple generations in globular clusters and predicted that 2G stars form centrally concentrated in the inner regions of the 1G system (see e.g. D’Ercole et al. 2008; Bekki 2011; Bekki et al. 2017; Calura et al. 2019; Lacchin et al. 2021, 2022; Yaghoobi et al. 2022a, b). While long-term dynamical evolution can gradually erase information of the initial conditions of these systems, some clusters can still preserve this information (see e.g. Vesperini et al. 2021; Livernois et al. 2024), as numerous observational studies demonstrate (Richer et al. 2013; Bellini et al. 2015, 2018; Milone et al. 2018; Dalessandro et al. 2019, 2024; Libralato et al. 2023; Leitinger et al. 2023; Cordoni et al. 2020, 2025; Mehta et al. 2025).
As the cluster evolves due to the effects of two-body relaxation, stellar interactions and energy exchange drive the system toward energy equipartition, a dynamical state where at a given distance from the cluster center stars with different masses have equal kinetic energy, $K=\frac{1}{2}m\sigma^{2}$, and thus a velocity dispersion, $\sigma$, increasing for smaller stellar masses (see e.g. Spitzer 1969). Theoretical and observational works suggest that GCs do not reach full energy equipartition, but are characterized by a state of partial equipartition (see e.g. Trenti & van der Marel 2013; Bianchini et al. 2016; Libralato et al. 2018; Watkins et al. 2022; Pavlík & Vesperini 2021, 2022; Aros & Vesperini 2023).
@@
However, measuring the degree of energy equipartition of the multiple populations in GCs is challenging. Generally, high-precision proper motion measurements in GCs are more commonly available for bright stars, such as those on the red giant branch (RGB), asymptotic giant branch (AGB), and horizontal branch (HB), which limits the analysis to a narrow mass range. This excludes the entire main sequence (MS), which encompasses a significantly broader range of stellar masses compared to giant stars. Furthermore, high-precision photometry with filters that effectively separate multiple populations is also required.
@@
In the past few years, thanks to progress in measuring kinematics of low-mass stars with the HST, the measurement of energy equipartition in GCs has become possible. Watkins et al. (2022) measured the degree of energy equipartition in the central regions of nine GCs by using photometric and astrometric data from the Hubble Space Telescope Proper Motion (HSTPROMO) catalogs of Galactic GCs, finding that no cluster is at or near full equipartition. Similarly, Libralato et al. (2022) found a partial state of energy equipartition in NGC 5904. Milone et al. (2025) explored the degree of energy equipartition in NGC 6397, finding that stars in the cluster exhibit only a very small degree of energy equipartition. Heyl et al. (2017) investigated mass-dependent kinematics in 47 Tucanae, finding a weak dependence of proper motions on stellar mass just beyond the cluster’s half-light radius, and a stronger dependence at the cluster center. In the context of multiple stellar populations, only one cluster has been studied, and within a small radial range. Bellini et al. (2018) measured the degree of energy equipartition for 1G and 2G stars in $\omega$ Centauri, finding that 2G stars present some degree of energy equipartition, while 1G stars are not in equipartition.
@@
Recently, we initiated a project to investigate the GC 47 Tucanae and its stellar populations using JWST observations from GO 2560 (PI: A. F. Marino) and archival data from major telescopes, including HST, JWST, and Gaia. Early results from this project include: i) the identification and chemical characterization of multiple stellar populations among very-low-mass stars using deep Near Infrared Camera (NIRCam) photometry (Milone et al. 2023b; Marino et al. 2024a; Legnardi et al. 2024); ii) the first detection of the brown dwarf sequence and the discovery of gaps and discontinuities within it (Marino et al. 2024a); iii) the inaugural spectroscopic study of multiple populations among M-dwarfs, based on Near-Infrared Spectrograph (NIRSpec) data (Marino et al. 2024b); and iv) the exploration of binary systems among first-generation and second-generation stars (Milone et al. 2025).
@@
In this paper, we continue our investigation into the multiple stellar populations of 47 Tucanae focusing on the kinematics of these populations. By leveraging a combination of datasets from JWST, HST, Gaia, and ground-based telescopes, we identify stellar populations across a mass range from $\sim$0.1 to 0.9 $M_{\odot}$ and derive precise proper motions for these stars. For the first time, we measure the differences in the degree of energy equipartition between 1G and 2G stars in 47 Tucanae, within an extensive radial range from approximately 1.5 to 4.5 $R_{h}$. We further analyze the internal kinematics of 1G and 2G stars, including their anisotropy profiles, over an extensive radial range that extends from the center of the cluster to over 30 arcminutes. This paper is organized as follows: Section 2 describes our photometric and astrometric catalogs; in Section 4, we outline our procedure for deriving stellar masses; the methodology and results regarding velocity dispersion and anisotropy profiles are discussed in Section 5; Section 5.6 presents the energy equipartition models and results; and Section 6 provides a discussion of the results, including energy equipartition radial profiles for 1G and 2G stars.
@@ 2 Data
In this section, we describe the dataset used to identify multiple stellar populations in the various fields of view analyzed in this study, illustrated in Figure 1. The cumulative exposure times across different years of observation are illustrated in the histogram in Figure 2. The data for fields A, B, C, D, and the inner field are available in the Mikulski Archive for Space Telescopes (MAST) [^1] Inner field: http://dx.doi.org/10.17909/gn0t-6945;
Field A: http://dx.doi.org/10.17909/vha2-8275;
Field B: http://dx.doi.org/10.17909/dyjj-z730;
Field C: http://dx.doi.org/10.17909/bcf5-c348;
Field D: http://dx.doi.org/10.17909/0abv-jg85.
@@
Field A, situated approximately 7 arcminutes from the cluster center, has been extensively observed with the HST, providing deep optical and near-infrared data. With this dataset, we were able to identify multiple populations among lower- ($M<0.45M_{\odot}$) and upper-MS ($0.780.67<M/M_{\odot}<$ 0.78) stars. This dataset was previously used by Milone et al. (2023b), which presents a detailed study of the chemical composition of multiple populations among very low-mass stars in 47 Tucanae.
Fields B and C, situated at $\sim 8.5$ and $\sim 11$ arcminutes from the cluster center, respectively, were observed with JWST/NIRCam as part of the program GO-2560 (PI: A. F. Marino). To classify populations in these fields, we combined NIRCam photometry in the F115W and F322W2 filters with observations taken with the HST Ultraviolet-VISible (UVIS) channel of the Wide Field Camera 3 (WFC3) in the F606W filter, and with NIR/WFC3 in the F110W and F160W filters. See Marino et al. (2024a) and Milone et al. (2023b) for a detailed description of these datasets and a study of multiple populations among very low-mass stars and brown dwarfs in 47 Tucanae.
@@
To measure stellar fluxes and positions in HST images, we followed a two-step procedure. First-step photometry for infrared images was performed using the FORTRAN code hst1pass (Anderson 2022) to find and measure bright and unsaturated sources in the individual exposures. For optical images, we generally used the img2xym computer program. Originally developed by Jay Anderson for HST image reduction (e.g., Anderson & King 2006), this approach measures stellar fluxes and positions in each image separately, using a spatially variable PSF model (Anderson & King 2000) along with a perturbation PSF to account for subtle focus variations. The perturbation PSF was derived from bright, unsaturated, and isolated stars. Photometric measurements from different exposures and detectors were standardized to a common zero-point by referencing the deepest exposure in each filter. This was achieved by using well-measured bright stars to compute the magnitude offsets between individual exposures and the reference frame, allowing all measurements to be transformed into a common master frame. Geometric distortion corrections were applied using the solutions from Milone et al. (2023b). Stellar positions were transformed into a common reference frame aligned with the Gaia DR3 catalog, ensuring that the x- and y-coordinates corresponded to west and north, respectively.
Second-step photometry was performed using the FORTRAN-based software KS2, developed by Jay Anderson as an advancement of the kitchen sinc algorithm (Anderson et al. 2008). KS2 implements three distinct techniques for stellar measurements. Method I, optimal for bright stars, determines magnitudes and positions by fitting the most suitable effective Point Spread Function (PSF) model (e.g., Anderson & King 2000). These values are computed individually for each image and subsequently averaged to refine the final measurements. Method II, designed to optimize astrometry and photometry for faint sources, integrates information from multiple exposures simultaneously. In this approach, fluxes are derived by subtracting neighboring stars and performing aperture photometry within a 5 $\times$ 5 pixel grid. Method III operates similarly to Method II but employs circular aperture photometry with a 0.75-pixel radius, making it particularly effective in crowded fields (see Sabbi et al. 2016; Bellini et al. 2017; Milone et al. 2023a, for details). Photometric calibration follows the procedures in Milone et al. (2023a) and utilizes the photometric zero-points available on the Space Telescope Science Institute (STScI) webpage for WFC/ACS, UVIS/WFC3, and NIR/WFC3222https://www.stsci.edu/hst/instrumentation/acs/data-analysis/zeropoints, https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration. Geometric distortion corrections are applied using the solutions from Anderson & King (2006) and Bellini et al. (2011) for WFC/ACS and UVIS/WFC3, while NIR/WFC3 corrections follow Anderson (2022). To ensure high-quality photometry and astrometry, we implemented the selection criteria outlined by (Milone et al. 2023a, see their Section 2.4), leveraging KS2 diagnostic parameters to identify well-isolated stars with minimal root mean scatter in position and magnitude.
Similarly, for JWST/NIRCam data reduction, we performed the same two-step procedure. First-step photometry was obtained using the img2xym computer program described above. Second-step photometry was performed using KS2 (see Milone et al. 2023b; Marino et al. 2024a, for details). The final photometry was calibrated to the VEGA system using the zero-points provided on the STScI webpage for NIRCam [^3] https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-performance/nircam-absolute-flux-calibration-and-zeropoints.
@@
To analyze the kinematics of more massive 1G and 2G stars, we used three catalogs: Cordoni et al. (2025), Mehta et al. (2025), and Lee (2022). Mehta et al. (2025) used Gaia DR3 XP spectra (Gaia Collaboration et al. 2023) to derive synthetic photometric filters that distinguish multiple populations in the RGB, HB, and AGB out to the tidal radius of 47 Tucanae. Using ground-based photometry from Stetson et al. (2019), Cordoni et al. (2025) applied the $\Delta_{C_{U,B,I}}$ vs. $\Delta_{B,I}$ chromosome map (ChM) to separate 1G and 2G RGB stars in 47 Tucanae, out to distances of about 20 arcminutes. Finally, Lee (2022) compiled large field-of-view ($\sim 1^{\circ}\times 1^{\circ}$) ground-based photometry in filters designed to maximize the separation of 1G and 2G stars. We matched this catalogs with Gaia DR3 proper motions to examine kinematic differences between 1G and 2G stars. Together, the catalogs from Lee (2022); Mehta et al. (2025); Cordoni et al. (2025) cover stars from about 5 arcminutes to the tidal radius of 47 Tucanae.
@@
Figure 3: Photometric diagrams demonstrating the photometry used to disentangle the multiple populations in different radial regions of 47 Tucanae. Panel a) shows the HST UV photometry used to identify populations in the RGB. In panel b), we show the CMD created using HST optical filters for Field A. Panel c) demonstrates the CMD created for the lower MS and substellar objects with JWST filters for Field C. Panel d) displays HST IR photometry for disentangling populations among the lower MS. In panel e) we show Ca-CN photometry by Lee (2022), which is effective in identifying populations among RGB and HB stars. Panel f) displays the diagram with ground-based UBVI photometry from Stetson et al. (2019). Finally, in panel g), we show the CMD created with Gaia XP spectro-photometry (Mehta et al. 2025).
@@
3 Proper motions
To assess relative proper motions for HST and JWST data, we employed a methodology detailed in several previous studies (e.g., Anderson & King 2003; Piotto et al. 2012; Libralato et al. 2022; Milone et al. 2023b), which focuses on comparing stellar positions across images captured at different epochs.
@@ 3 Proper motions
To transform star coordinates from each catalog into this master frame, we utilized six-parameter linear transformations. The transformation process relied on bright, unsaturated cluster stars. This selection was conducted in two phases: initially, we identified likely members of 47 Tucanae based solely on their positions in the CMD, which allowed us to derive preliminary proper motions. In the second phase, we refined our criteria by examining each star’s position within both the CMD and the proper motion diagram, leading to a more accurate identification of cluster members. These members were then used to calculate improved proper motions. To further reduce uncertainties in our measurements, we adjusted stellar coordinates at each epoch by accounting for displacements due to proper motions. These corrected coordinates were then utilized to calculate improved transformations, further refining our estimates of proper motion.
To enhance alignment and minimize small residual distortions, accounting for CTE and geometric distortion residuals in the HST (ACS and UVIS) images, we implemented local transformations based on the nearest N reference stars, deliberately excluding target stars themselves from influencing these corrections (as described in Bellini et al. 2014). The selection of N and the corresponding magnitude range was optimized to balance capturing local variations while maintaining a statistically significant sample of reference stars. For stars near the extremes of the magnitude range, either very bright or very faint targets, we adjusted the magnitude selection criteria by using reference stars within the brightest or faintest magnitude bins instead of applying a fixed magnitude difference. This adaptive approach yielded a larger statistical sample and improved the correction process in regions where suitable reference stars were otherwise scarce.
We then plotted the x and y coordinates of each star against the observation epochs, fitting a weighted least-squares linear regression to these data points. The slope of this fitted line, along with its uncertainty, yields our best estimate of the star’s proper motion.